OpenStudy (anonymous):

can someone PLEASE help me on some Algebra 1 homework project questions?? I'd be soooo grateful!! + medal

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

thank you!!!! i'll screenshot the questions :D

OpenStudy (anonymous):

@alysha_32 even if you only know how to do one or 2 questions it'd really help! :)

OpenStudy (anonymous):

Ben= f(x)=2x^2+2

OpenStudy (anonymous):

awesome thanks. do you know how to find #2? @alysha_32

OpenStudy (lynfran):

@alysha_32 are u sure thats correct?

OpenStudy (anonymous):

oh is it not?... would you care to help me if you know how please @LynFran

OpenStudy (anonymous):

i thought it is cuz 2(2)^2+2= 18

OpenStudy (anonymous):

and 2(1)^2+2=6

OpenStudy (lynfran):

2(2)^2+2=? 2(4)+2=? 8+2=10.... 2(1)^2+2=? 2(1)+2=? 2+2=4.... Any way try this... let the # of days=n so \[4(n)^{2}+3-1\] let test it... \[4(0)^{2}+3-1=?\]\[4(0)+2=2\] \[4(1)^{2}+3-1\]\[4(1)+2\]\[4+2=6\] \[4(2)^{2}+3-1\]\[4(4)+2\]\[16+2=18\]

OpenStudy (anonymous):

are you using PEMDAS? @LynFran ?

OpenStudy (anonymous):

@LynFran I'm confused which question was that for?

OpenStudy (anonymous):

#1? ^

OpenStudy (lynfran):

yes @alysha_32 and @sparkles456

OpenStudy (anonymous):

how would I give the answer to #1 then to sum all of that up @LynFran & sorry if you think i sound clueless

OpenStudy (lynfran):

f(x)=4(x)^2 +3-1

OpenStudy (anonymous):

hmmm ok I appreciate the help & would you mind helping me on a few more of those? i could give you a medal if you do and keep in mind I'm new to open study aswell so I'm not sure if i could reward you in another way too @LynFran

OpenStudy (anonymous):

what i don't understand is how you both got completely different answers lol @alysha_32 @LynFran if it helps the lesson is with exponential functions

OpenStudy (lynfran):